# Heuristic Learning with Graph Neural Networks: A Unified Framework for Link Prediction

### v1

@@ Abstract.
Link prediction is a fundamental task in graph learning, inherently shaped by the topology of the graph. While traditional heuristics are grounded in graph topology, they encounter challenges in generalizing across diverse graphs. Recent research efforts have aimed to leverage the potential of heuristics, yet a unified formulation accommodating both local and global heuristics remains undiscovered. Drawing insights from the fact that both local and global heuristics can be represented by adjacency matrix multiplications, we propose a unified matrix formulation to accommodate and generalize various heuristics. We further propose the Heuristic Learning Graph Neural Network (HL-GNN) to efficiently implement the formulation. HL-GNN adopts intra-layer propagation and inter-layer connections, allowing it to reach a depth of around 20 layers with lower time complexity than GCN. HL-GNN is proven to be more expressive than heuristics and conventional GNNs, and it can adaptively trade-off between node features and topological information. Extensive experiments on the Planetoid, Amazon, and OGB datasets underscore the effectiveness and efficiency of HL-GNN. It outperforms existing methods by a large margin in prediction performance. Additionally, HL-GNN is several orders of magnitude faster than heuristic-inspired methods while requiring only a few trainable parameters. The case study further demonstrates that the generalized heuristics and learned weights are highly interpretable. The code is available at https://github.com/LARS-research/HL-GNN [^1] Correspondence is to Q. Yao at qyaoaa@tsinghua.edu.cn.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 25–29, 2024; Barcelona, Spain.booktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spainisbn: 979-8-4007-0490-1/24/08doi: 10.1145/XXXXXX.XXXXXXccs: Theory of computation Graph algorithms analysisccs: Computing methodologies Neural networks
@@ 1. Introduction
While heuristics prove effective in link prediction tasks, they inherently capture specific topology patterns, posing challenges in their generalization to diverse graphs (Zhu et al., 2021; Yun et al., 2021). Moreover, heuristics are unable to leverage node features, limiting their expressiveness and efficacy on attributed graphs (Zhang and Chen, 2018). To make heuristics more universal and general, recent research efforts have been directed toward establishing formulations for heuristics and learning heuristics from these formulations. Notable examples include SEAL (Zhang and Chen, 2018), NBFNet (Zhu et al., 2021), and Neo-GNN (Yun et al., 2021). SEAL’s $\gamma$-decaying framework and NBFNet’s path formulation are tailored for global heuristics, while Neo-GNN’s MLP framework is tailored for local ones. To obtain multi-range topological information, a unified formulation that accommodates both local and global heuristics is necessary, yet it remains undiscovered.
@@
To learn generalized heuristics and acquire multi-range information, we propose the Heuristic Learning Graph Neural Network (HL-GNN) to efficiently implement the formulation. HL-GNN incorporates intra-layer propagation and inter-layer connections while excluding transformation and activation functions. This design enables HL-GNN to effectively reach a depth of around 20 layers, while only requiring the training of a global GNN with a time complexity even lower than GCN. We derive the theoretical properties of the polynomial graph filter, which governs the trade-off between node features and topological information. Moreover, we prove that the unified formulation is strictly more expressive than individual heuristics, and HL-GNN is strictly more expressive than both individual heuristics and conventional GNNs.
@@
  • We propose HL-GNN to efficiently implement the formulation, capable of reaching a depth of around 20 layers with lower time complexity than GCN. HL-GNN is strictly more expressive than both individual heuristics and conventional GNNs.
@@ 2.2. Link Prediction
Recent advancements have focused on enhancing GNNs with valuable topological information. Subgraph GNNs like SEAL (Zhang and Chen, 2018), GraIL (Teru et al., 2020), and SUREL (Yin et al., 2022) explicitly encode subgraphs around node pairs. However, a drawback of these methods is that they require the running of a subgraph GNN with the labeling trick for each link during training and inference. Taking a different perspective, models like NBFNet (Zhu et al., 2021) and RED-GNN (Zhang and Yao, 2022) adopt source-specific message passing, drawing inspiration from global heuristics. While offering advantages in terms of capturing broader topological information, they require training a global GNN for each source node. Some methods opt for a single global GNN to improve scalability and efficiency. Neo-GNN (Yun et al., 2021) uses two MLPs, while SEG (Ai et al., 2022) uses a GCN layer and an MLP to approximate a heuristic function. BUDDY (Chamberlain et al., 2022) develops a novel GNN that passes subgraph sketches as messages, eliminating the need for subgraph construction. However, these methods primarily focus on local topological information and struggle to capture global topological information. In contrast, the proposed HL-GNN can capture long-range information up to 20 hops while only requiring the training of a global GNN. Further details about the comparison of HL-GNN with existing methods are provided in Section 4.2.
@@ 3. Unified Heuristic Formulation
(1) $\boldsymbol{H}=\sum\nolimits_{l=0}^{L}\left(\beta^{(l)}\prod\nolimits_{m=0}^{l}\mathbb{A}^{(m)}\right),$
@@
Existing heuristics are primarily handcrafted and may not be optimal for real-world graphs. Leveraging the propagation operators, weight parameters and maximum order offers the potential to learn generalized, possibly more effective heuristics, which we will discuss in Section 5.5.1 and Appendix LABEL:para:heuristic_advance.
@@ 4.1.2. Architecture
(2) $\boldsymbol{Z}^{(0)}=\boldsymbol{X},\quad\boldsymbol{Z}^{(l)}=\mathbb{A}^{(l)}\boldsymbol{Z}^{(l-1)},\quad\boldsymbol{Z}=\sum\nolimits_{l=0}^{L}\beta^{(l)}\boldsymbol{Z}^{(l)},$
@@
We assert the relationship between the learned representations $\boldsymbol{Z}$ and the heuristic formulation $\boldsymbol{H}$ in Proposition 4.1. The proof for Proposition 4.1 can be found in Appendix A.2.
@@ 4.2. Comparison with Existing Methods
HL-GNN has a time complexity of $\mathcal{O}(LMF)$, which is the lowest among the compared methods. Unlike conventional GNNs, HL-GNN solely utilizes propagation mechanisms and omits transformation and activation functions. SEAL requires running a subgraph GNN with the labeling trick for each link, and NBFNet requires running a global GNN for each source node during training and inference. In contrast, HL-GNN only requires running a single global GNN during training and inference. Furthermore, HL-GNN avoids the need to extract topological information from common neighbors and subgraph sketches, as required by Neo-GNN and BUDDY, respectively.
@@ 5.1.3. Experimental settings
We include a linear layer as preprocessing before HL-GNN to align the dimension of node features with the hidden channels of HL-GNN. We also leverage node embeddings on the OGB datasets to enhance the expressiveness of node representations. For the ogbl-collab dataset, we follow OGB’s guidelines and use the validation set for training. We evaluate HL-GNN over 10 runs without fixing the random seed. More details about the experiments are provided in Appendix B.2.
@@ 5.2. Main Results
As shown in Table 4, HL-GNN consistently outperforms all the baselines on all of the datasets, highlighting its effectiveness and robustness for link prediction tasks. Table 4 reports the averaged results with standard deviations. Notably, HL-GNN achieves a remarkable gain of 7.0% and 16.7% in Hits@100 compared to the second-best method on the Planetoid datasets Cora and Pubmed, respectively. Moreover, our HL-GNN demonstrates its ability to handle large-scale graphs effectively, as evidenced by its superior performance on the OGB datasets. Specifically, HL-GNN achieves a gain of 13.9% in Hits@100 on ogbl-ppa, and achieves 68.11% Hits@50 on ogbl-collab and 89.43% MRR on ogbl-citation2. Even when node features are absent or of low quality, HL-GNN maintains consistent performance by learning embeddings for each node, as demonstrated on datasets like ogbl-ddi, which lack node attributes.
HL-GNN outperforms all listed heuristics, indicating its capacity to generalize heuristics and integrate them with node features. According to Table 4, local heuristics like CN and RA perform better than global heuristics on the OGB datasets, while global heuristics like KI and RWR perform better on the Planetoid and Amazon datasets. This underscores the importance of establishing a unified formulation that accommodates both local and global heuristics. Notably, we can use the configuration in Table 2 to recover the heuristic RA from HL-GNN without training, achieving a performance of 49.33% Hits@100 on ogbl-ppa. This result serves as a compelling lower bound for HL-GNN’s performance on ogbl-ppa. HL-GNN significantly outperforms conventional GNNs like GCN and GAT across all datasets, attributed to its higher expressiveness than conventional GNNs. HL-GNN also surpasses existing heuristic-inspired GNN methods, including SEAL, NBFNet, Neo-GNN, and BUDDY, suggesting that integrating information from multiple ranges is beneficial for link prediction tasks.
@@ 5.3.1. Different information ranges
(b) Citeseer.
Figure 2. Ablation study on information ranges. We compare HL-GNN with two GNN variants, each focusing on either local or global topological information, using different GNN depths $L$ across various datasets.
@@ 5.3.2. Sufficient model depths
Figure 3. Test performance of link prediction on the Planetoid and OGB datasets with varying GNN depths $L$.
@@ 5.5.1. Generalized heuristics and learned weights
Leveraging the capabilities of the unified formulation, we can derive generalized heuristics by analyzing the learned parameters of HL-GNN. The generalized heuristics and learned weights $\beta^{(l)}$ provide insights into the graph-structured data. The learned weights $\beta^{(l)}$ are visually depicted in Figure 4.
For the Cora and Citeseer datasets, the learned weights monotonically decrease, indicating that the graph filter serves as a low-pass filter. The weight $\beta^{(0)}$ has the largest magnitude, suggesting that crucial information is primarily contained in node features. Local topological information from nearby neighbors plays a major role, while global topological information from distant nodes serves as a complementary factor. Conversely, for the ogbl-collab and ogbl-ddi datasets, the weights do not monotonically increase or decrease. Instead, they experience a significant change, especially in the first 5 layers, indicating that the graph filter serves as a high-pass filter. The weight $\beta^{(2)}$ has the largest magnitude, suggesting that crucial information lies in local topology rather than node features. Moreover, for large values of $l$ on the ogbl-collab and ogbl-ddi datasets, the weights $\beta^{(l)}$ become negative, suggesting that global topological information from distant nodes compensates for excessive information from nearby neighbors. According to Figure 4, the learnable weights $\beta^{(l)}$ adaptively determine the properties of the graph filter, perfectly aligning with Theorem LABEL:prop:graph_filter. These weights govern the trade-off between node features and topological information, enabling the adaptive integration of multi-range topological information.
@@ 5.5.2. Leveraging generalized heuristics
With the generalized heuristic for each dataset, there is no need to train a GNN and an MLP predictor from scratch. Instead, we can simply follow the generalized heuristic and train an MLP predictor only, which is significantly more efficient than training from scratch. The performance of training the MLP alone is comparable to training from scratch, but it converges more quickly. The training time for training from scratch versus training only the MLP is shown in Table 7. The slight decrease in performance can likely be attributed to the fact that, when training the GNN and MLP together, the gradients flow through both blocks, allowing them to adapt to each other. In contrast, training the MLP alone limits its ability to capture complex interactions between the two blocks.
@@ 6. Conclusion
We introduce a unified formulation that accommodates and generalizes both local and global heuristics using propagation operators and weight parameters. Additionally, we propose HL-GNN, which efficiently implements this formulation. HL-GNN combines intra-layer propagation and inter-layer connections, allowing the integration of multi-range information. Experiments demonstrate that HL-GNN achieves state-of-the-art performance and efficiency. This study is confined to undirected graphs; for directed graphs, we preprocess them by converting them into undirected graphs. Extensions to multi-relational graphs, such as knowledge graphs, are left for future work.
@@ Acknowledgment
Q. Yao was in part supported by National Key Research and Development Program of China under Grant 2023YFB2903904 and NSFC (No. 92270106).
@@ Appendix A Proofs
Building upon above lemmas, it becomes evident that a diverse range of local and global heuristics can be represented in matrix form. Consequently, our formulation is capable of accommodating a wide range of local and global heuristics. This observation serves to prove Proposition 3.3. ∎
Table 8. Statistics of datasets used in experiments.
Cora Citeseer Pubmed Photo Computers collab ddi
# Nodes 2,708 3,327 18,717 7,650 13,752 235,868 4,267
# Edges 5,278 4,676 44,327 238,162 491,722 1,285,465 1,334,889
# Features 1,433 3,703 500 745 767 128 /
Dataset Split random random random random random fixed fixed
Average Degree 3.90 2.81 4.73 62.26 71.51 5.45 312.84
@@ B.1. Detailed Information about Datasets
We employ a diverse set of benchmark datasets to comprehensively evaluate the link prediction performance of HL-GNN. Each dataset captures distinct aspects of relationships and presents unique challenges for link prediction. We utilize seven datasets from three sources: Planetoid (Sen et al., 2008), Amazon (Shchur et al., 2018), and OGB (Hu et al., 2020). The Planetoid datasets include Cora, Citeseer, and Pubmed. The Amazon datasets include Photo and Computers. The OGB datasets include ogbl-collab and ogbl-ddi. The statistics of each dataset are shown in Table 8.

### v2

@@ Abstract.
Link prediction is a fundamental task in graph learning, inherently shaped by the topology of the graph. While traditional heuristics are grounded in graph topology, they encounter challenges in generalizing across diverse graphs. Recent research efforts have aimed to leverage the potential of heuristics, yet a unified formulation accommodating both local and global heuristics remains undiscovered. Drawing insights from the fact that both local and global heuristics can be represented by adjacency matrix multiplications, we propose a unified matrix formulation to accommodate and generalize various heuristics. We further propose the Heuristic Learning Graph Neural Network (HL-GNN) to efficiently implement the formulation. HL-GNN adopts intra-layer propagation and inter-layer connections, allowing it to reach a depth of around 20 layers with lower time complexity than GCN. Extensive experiments on the Planetoid, Amazon, and OGB datasets underscore the effectiveness and efficiency of HL-GNN. It outperforms existing methods by a large margin in prediction performance. Additionally, HL-GNN is several orders of magnitude faster than heuristic-inspired methods while requiring only a few trainable parameters. The case study further demonstrates that the generalized heuristics and learned weights are highly interpretable. The code is available at https://github.com/LARS-research/HL-GNN.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 25–29, 2024; Barcelona, Spain.booktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671946isbn: 979-8-4007-0490-1/24/08ccs: Theory of computation Graph algorithms analysisccs: Computing methodologies Neural networks
@@ 1. Introduction
While heuristics prove effective in link prediction tasks, they inherently capture specific topology patterns, posing challenges in their generalization to diverse graphs (Zhu et al., 2021; Yun et al., 2021). Moreover, heuristics are unable to leverage node features, limiting their efficacy on attributed graphs (Zhang and Chen, 2018). To make heuristics more universal and general, recent research efforts have been directed toward establishing formulations for heuristics and learning heuristics from these formulations. Notable examples include SEAL (Zhang and Chen, 2018), NBFNet (Zhu et al., 2021), and Neo-GNN (Yun et al., 2021). SEAL’s $\gamma$-decaying framework and NBFNet’s path formulation are tailored for global heuristics, while Neo-GNN’s MLP framework is tailored for local ones. To obtain multi-range topological information, a unified formulation that accommodates both local and global heuristics is necessary, yet it remains undiscovered.
@@
To learn generalized heuristics and acquire multi-range information, we propose the Heuristic Learning Graph Neural Network (HL-GNN) to efficiently implement the formulation. HL-GNN incorporates intra-layer propagation and inter-layer connections while excluding transformation and activation functions. This enables HL-GNN to effectively reach a depth of around 20 layers, while only requiring the training of a global GNN with a time complexity even lower than GCN. The adaptive weights in HL-GNN facilitate the integration of multi-range topological information, and govern the trade-off between node features and topological information.
@@
  • We propose HL-GNN to efficiently implement the formulation, capable of reaching a depth of around 20 layers with lower time complexity than GCN. HL-GNN can adaptively balance the trade-off between node features and topological information.
@@ 2.2. Link Prediction
Recent advancements have focused on enhancing GNNs with valuable topological information. Subgraph GNNs like SEAL (Zhang and Chen, 2018), GraIL (Teru et al., 2020), and SUREL (Yin et al., 2022) explicitly encode subgraphs around node pairs. However, they require the running of a subgraph GNN with the labeling trick for each link during training and inference. Taking a different perspective, models like NBFNet (Zhu et al., 2021) and RED-GNN (Zhang and Yao, 2022) adopt source-specific message passing, drawing inspiration from global heuristics. However, they require training a global GNN for each source node. Some methods opt for a single global GNN to improve scalability and efficiency. Neo-GNN (Yun et al., 2021) uses two MLPs, while SEG (Ai et al., 2022) uses a GCN layer and an MLP to approximate a heuristic function. BUDDY (Chamberlain et al., 2022) develops a novel GNN that passes subgraph sketches as messages. However, these methods primarily focus on local topological information and struggle to capture global topological information. In contrast, the proposed HL-GNN can capture long-range information up to 20 hops while only requiring the training of a global GNN. Further details about the comparison of HL-GNN with existing methods are provided in Section 4.2.
@@ 3. Unified Heuristic Formulation
(1) $\boldsymbol{H}=\sum_{l=0}^{L}\left(\beta^{(l)}\prod_{m=0}^{l}\mathbb{A}^{(m)}\right),$
@@
Existing heuristics are primarily handcrafted and may not be optimal for real-world graphs. Leveraging the propagation operators, weight parameters and maximum order offers the potential to learn generalized, possibly more effective heuristics, which we will discuss in Section 5.5.1 and Appendix D.
@@ 4.1.2. Architecture
(2) $\boldsymbol{Z}^{(0)}=\boldsymbol{X},\quad\boldsymbol{Z}^{(l)}=\mathbb{A}^{(l)}\boldsymbol{Z}^{(l-1)},\quad\boldsymbol{Z}=\sum_{l=0}^{L}\beta^{(l)}\boldsymbol{Z}^{(l)},$
@@
HL-GNN employs intra-layer propagation and inter-layer connections as described in Equation (2). The salient trait is its elimination of representation transformation and non-linear activation at each layer, requiring only a few trainable parameters. We assert the relationship between the learned representations $\boldsymbol{Z}$ and the heuristic formulation $\boldsymbol{H}$ in Proposition 4.1. The proof for Proposition 4.1 can be found in Appendix A.2.
@@ 4.2. Comparison with Existing Methods
HL-GNN has a time complexity of $\mathcal{O}(LMF)$, which is the lowest among the compared methods. Unlike conventional GNNs, HL-GNN solely utilizes propagation mechanisms and omits transformation and activation functions. SEAL requires running a subgraph GNN with the labeling trick for each link, and NBFNet requires running a global GNN for each source node during training and inference. In contrast, HL-GNN only requires running a single global GNN during training and inference. Furthermore, HL-GNN avoids the need to extract topological information from common neighbors and subgraph sketches, as required by Neo-GNN and BUDDY, respectively. A detailed time complexity analysis of each method is included in Appendix E.
@@ 5.1.3. Experimental settings
We include a linear layer as preprocessing before HL-GNN to align the dimension of node features with the hidden channels of HL-GNN. We also leverage node embeddings on the OGB datasets to enhance the node representations. For the ogbl-collab dataset, we follow OGB’s guidelines and use the validation set for training. We evaluate HL-GNN over 10 runs without fixing the random seed. More details about the experiments are provided in Appendix B.2.
@@ 5.2. Main Results
As shown in Table 4, HL-GNN consistently outperforms all the baselines on all of the datasets, highlighting its effectiveness and robustness for link prediction tasks. Table 4 reports the averaged results with standard deviations. Notably, HL-GNN achieves a remarkable gain of 7.0% and 16.7% in Hits@100 compared to the second-best method on the Planetoid datasets Cora and Pubmed, respectively. Moreover, our HL-GNN demonstrates its ability to handle large-scale graphs effectively, as evidenced by its superior performance on the OGB datasets. Specifically, HL-GNN achieves a gain of 13.9% in Hits@100 on ogbl-ppa, and achieves 68.11% Hits@50 on ogbl-collab and 89.43% MRR on ogbl-citation2. Even when node features are absent or of low quality, HL-GNN maintains consistent performance by learning embeddings for each node, as demonstrated on datasets like ogbl-ddi, which lack node features.
HL-GNN outperforms all listed heuristics, indicating its capacity to generalize heuristics and integrate them with node features. According to Table 4, local heuristics like CN and RA perform better than global heuristics on the OGB datasets, while global heuristics like KI and RWR perform better on the Planetoid and Amazon datasets. This underscores the importance of establishing a unified formulation that accommodates both local and global heuristics. Notably, we can use the configuration in Table 2 to recover the heuristic RA from HL-GNN without training, achieving a performance of 49.33% Hits@100 on ogbl-ppa. This result serves as a compelling lower bound for HL-GNN’s performance on ogbl-ppa. HL-GNN significantly outperforms conventional GNNs like GCN and GAT across all datasets. Additionally, HL-GNN also surpasses existing heuristic-inspired GNN methods, including SEAL, NBFNet, Neo-GNN, and BUDDY, suggesting that integrating information from multiple ranges is beneficial for link prediction tasks.
@@ 5.3.1. Different information ranges
(b) ogbl-collab.
Figure 2. Ablation study on information ranges. We compare HL-GNN with two GNN variants, focusing on either local or global topological information, with different GNN depths.
@@ 5.3.2. Sufficient model depths
Figure 3. Test performance on the Planetoid and OGB datasets with different GNN depths.
@@ 5.5.1. Generalized heuristics and learned weights
Leveraging the capabilities of the unified formulation, we can derive generalized heuristics by analyzing the learned parameters of HL-GNN. The generalized heuristics and learned weights $\beta^{(l)}$ provide insights into the graph-structured data. The learned weights $\beta^{(l)}$ are visually depicted in Figure 4. Due to space constraints, the formulas for the generalized heuristics are provided in Appendix C.2.
For the Cora and Citeseer datasets, the learned weights monotonically decrease, indicating that the graph filter serves as a low-pass filter. The weight $\beta^{(0)}$ has the largest magnitude, suggesting that crucial information is primarily contained in node features. Local topological information from nearby neighbors plays a major role, while global topological information from distant nodes serves as a complementary factor. Conversely, for the ogbl-collab and ogbl-ddi datasets, the weights do not monotonically increase or decrease. Instead, they experience a significant change, especially in the first 5 layers, indicating that the graph filter serves as a high-pass filter. The weight $\beta^{(2)}$ has the largest magnitude, suggesting that crucial information lies in local topology rather than node features. Moreover, for large values of $l$ on the ogbl-collab and ogbl-ddi datasets, the weights $\beta^{(l)}$ become negative, suggesting that global topological information from distant nodes compensates for excessive information from nearby neighbors. The learnable weights $\beta^{(l)}$ govern the trade-off between node features and topological information, enabling the adaptive integration of multi-range topological information.
@@ 5.5.2. Leveraging generalized heuristics
With the generalized heuristic for each dataset, there is no need to train a GNN and an MLP predictor from scratch. Instead, we can simply follow the generalized heuristic and train an MLP predictor only, which is significantly more efficient than training from scratch. The performance of training the MLP alone is comparable to training from scratch, but it converges more quickly. The training time for training from scratch versus training only the MLP is shown in Table 7, while the performance details are provided in Appendix C.3. The slight decrease in performance can likely be attributed to the fact that, when training the GNN and MLP together, the gradients flow through both blocks, allowing them to adapt to each other. In contrast, training the MLP alone limits its ability to capture complex interactions between the two blocks.
@@ 6. Conclusion
We introduce a unified formulation that accommodates and generalizes both local and global heuristics using propagation operators and weight parameters. Additionally, we propose HL-GNN, which efficiently implements this formulation. HL-GNN combines intra-layer propagation and inter-layer connections, allowing the integration of multi-range topological information. Experiments demonstrate that HL-GNN achieves state-of-the-art performance and efficiency. This study is confined to undirected graphs; for directed graphs, we preprocess them by converting them into undirected graphs. Extensions to multi-relational graphs, such as knowledge graphs, are left for future work.
@@ Acknowledgment
This work is supported by National Key Research and Development Program of China (under Grant No.2023YFB2903904), National Natural Science Foundation of China (under Grant No.92270106) and Beijing Natural Science Foundation (under Grant No.4242039).
@@ Appendix A Proofs
We provide detailed derivations of the matrix forms of heuristics and demonstrate their alignment with the unified heuristic formulation under specific configurations.
@@
Building upon above lemmas, it becomes evident that a diverse range of local and global heuristics can be represented in matrix form. Our formulation is capable of accommodating a wide range of local and global heuristics under specific configurations. This serves to prove Proposition 3.3. ∎
@@ B.1. Detailed Information about Datasets
Table 8. Statistics of the Planetoid, Amazon, and OGB datasets used in the experiments.
Cora Citeseer Pubmed Photo Computers collab ddi ppa citation2
# Nodes 2,708 3,327 18,717 7,650 13,752 235,868 4,267 576,289 2,927,963
# Edges 5,278 4,676 44,327 238,162 491,722 1,285,465 1,334,889 30,326,273 30,561,187
# Features 1,433 3,703 500 745 767 128 / 58 128
Dataset Split random random random random random time time throughput protein
Average Degree 3.90 2.81 4.73 62.26 71.51 5.45 312.84 52.62 10.44
We employ a diverse set of benchmark datasets to comprehensively evaluate the link prediction performance of HL-GNN. Each dataset captures distinct aspects of relationships and presents unique challenges for link prediction. We utilize nine datasets from three sources: Planetoid (Sen et al., 2008), Amazon (Shchur et al., 2018), and OGB (Hu et al., 2020). The Planetoid datasets include Cora, Citeseer, and Pubmed. The Amazon datasets include Photo and Computers. The OGB datasets include ogbl-collab, ogbl-ddi, ogbl-ppa, and ogbl-citation2. The statistics of each dataset are shown in Table 8.
@@
  • ogbl-ppa: This dataset represents a protein-protein association network. Nodes correspond to proteins, and edges indicate associations based on various biological evidence. Node features include protein sequences and structural information.
  • ogbl-citation2: This dataset is a citation network that focuses on the dynamics of citations. Nodes represent papers, and edges indicate citation relationships. Node features consist of paper metadata.